# Collective Outlier Detection and Enumeration with Conformalized Closed Testing

### v2

@@ 2.6 Data-Driven Tuning
We extend Algorithm 2 to leverage a potentially diverse suite of classifiers and local testing procedures. The goal is to approximately maximize power while controlling the type-I errors. Our solution, outlined by Algorithm 3, is inspired by similar approaches proposed by Liang et al. (2024) and Marandon et al. (2024) in the context of individual outlier identification under FDR control. The idea is to embed the data-driven tuning step within an additional layer of sample splitting. As detailed below, this is crucial for maintaining the validity of local tests after selecting the most effective classifier and local testing method.
@@
6             Compute $d_{k,l}(S)$ by applying Algorithm 2 based on $\mathcal{A}^{(k)}$ and $\phi^{(l)}$, using $\mathcal{D}^{\text{train}}$ for training, $\mathcal{D}^{\text{tune}}$ for calibration, and $\tilde{\mathcal{D}}^{\text{test}}$ as the test set.
7             Compute $\hat{p}_{k,l}(S)$, the p-value for $H_{S}$ based on $\mathcal{A}^{(k)}$ and $\phi^{(l)}$.
@@
9
10Find $(\hat{k}(S),\hat{l}(S))=\arg\min_{k\in[K],l\in[L]}\arg\max_{k\in[K],l\in[L]}d_{k,l}(S)$; if not unique, look at $\hat{p}_{k,l}(S)$.
11
12Compute $d(S)$ by applying Algorithm 2 based on $\mathcal{A}^{(\hat{k}(S))}$ and $\phi^{(\hat{l}(S))}$, using $\mathcal{D}^{\text{train}}\cup\mathcal{D}^{\text{tune}}$ for training, $\mathcal{D}^{\text{cal}}$ for calibration, and $\mathcal{D}^{\text{test}}$ as the test set.
Output : A $(1-\alpha)$ lower bound $d(S)$ for the number of outliers in any $S\subseteq[n]$.
13
@@
The initial tuning module of Algorithm 3 sees the calibration and test data only through the lenses of the unordered collection $\{Z_{1},\ldots,Z_{m}\}\cup\{Z_{m+1},\ldots,Z_{m+n}\}$. Consequently, when Algorithm 2 is applied again in the inference step of Algorithm 3, using $(Z_{1},\ldots,Z_{m})$ as a calibration set and $(Z_{m+1},\ldots,Z_{m+n})$ as a test set, the scores $(Z_{1},\ldots,Z_{m+n})$ corresponding to inliers are still exchangeable conditional on the selected classifier and testing procedure. Thus, Algorithm 3 can often achieve high power without selection bias, as previewed in Figure 1 and confirmed in Section 3. Formally, Algorithm 3 also outputs a simultaneous lower bound $d(S)$ guaranteed to satisfy (9), as long as the local tests are valid.
@@
Algorithm 3 changes this picture a little. As we condition on the selected classifier and closed testing procedure, the scores of the inliers among $(Z_{1},\ldots,Z_{m+n})$, although still exchangeable, can no longer be independent, irrespective of whether they were initially obtained through one-class or binary classification. This is because such conditioning is similar to (but weaker than) conditioning on the unordered collection $\{Z_{1},\ldots,Z_{m}\}\cup\{Z_{m+1},\ldots,Z_{m+n}\}$, an operation that transforms the distribution of $(Z_{1},\ldots,Z_{m+n})$ into a finite-population distribution. However, this issue does not seem to have noticeable practical implications. This robustness can be explained by noting that, among the local testing procedures considered in this paper, independence is only needed to derive the large-sample asymptotic distribution of the WMW and Shirashi test statistics. (In small samples, their critical values can be obtained via permutations, under the sole assumption of exchangeability.) In large samples, however, conditioning on $\{Z_{1},\ldots,Z_{m}\}\cup\{Z_{m+1},\ldots,Z_{m+n}\}$ makes little practical difference because the cardinality of this set is large, since all our scores always take distinct values in practice.
@@ 3 Empirical Demonstrations
Section 3.1 applies ACODE to synthetic data, and Section 3.2 to the 2020 LHCO data (Kasieczka et al., 2021). Additional experiments are described in Appendix A4.
@@ 3.1.1 Setup
We apply ACODE on synthetic data from a distribution inspired by Liang et al. (2024) and Bates et al. (2023). Each observation $Z_{i}\in\mathbb{R}^{1000}$ is sampled from a multivariate Gaussian mixture $P_{Z}^{a}$, such that $Z_{i}=\sqrt{a}\,V_{i}+W_{i}$, for some constant $a\geq 0$ and vectors $V_{i},W_{i}\in\mathbb{R}^{1000}$. The inliers correspond to $a=1$ and the outliers to $a=0.7$. The elements of $V_{i}$ are standard Gaussian, while each element of $W_{i}$ is independent and uniformly distributed on a discrete set $\mathcal{W}\subseteq\mathbb{R}^{1000}$ with $|\mathcal{W}|=1000$. The vectors in $\mathcal{W}$ are independently sampled from Uniform($[-3,3]^{1000}$) prior to the first experiment. The numbers of inliers in the training, calibration, and tuning sets utilized by ACODE are 1000, 750, and 250, respectively.
@@
The testing procedures leveraged by ACODE are: Simes’ test with and without Storey’s correction, Fisher’s combination method, the WMW test, and the LMPI test described in Section 2.3.2, applied using $G(F)=F^{3}$. Recall that this choice of $G$ is optimal against Lehmann’s alternative with $k=3$, which may or may not fit our data well. An implementation of the LMPI test based on a data-driven estimate of $G$ will be considered later.
@@ 3.1.3 Selective Outlier Enumeration
Figure 3 reports on related experiments in which the goal is to construct a 90% lower confidence bounds for the number of outliers within a data-driven subset of test points, selected as those with the largest conformity scores. To facilitate the interpretation of these experiments, ACODE is applied using a fixed classification algorithm (a one-class support vector machine), not a suite of 6 different algorithms. This ensures the test subsets are always selected based on comparable conformity scores in all repetitions of the experiments. These results further demonstrate how the efficacy of different local testing procedures can vary in different situations. Simes’ method performs better with very small selected sets, while the LMPI test with $G(F)=F^{3}$ excels in cases with moderately large selected sets. Once more, ACODE can approximately maximize power by autonomously identifying the most effective classifier and local testing procedure for each scenario.
@@ 3.1.4 Hunting for Adversarial Anomalies
We now apply ACODE leveraging an additional local testing procedure, utilizing a more flexible implementation of the LMPI approach outlined in Section 2.3.2. This approach enables our method to even detect elusive anomalies hidden by an adversary.
@@
Without prior knowledge of the distribution of outlier scores, ACODE is applied as in previous experiments, using the same toolbox of six machine learning classifiers, but including a sixth local testing procedure. This procedure is a practical approximation of the LMPI test described in Section 2.3.2, using an empirical estimate $\hat{G}$ of $G$. This estimate is obtained by fitting a mixture of Beta distributions, while preserving the exchangeability between calibration and test scores, as detailed in Appendix A2.
The results reported in Figure 4 demonstrate that our method effectively detects these elusive adversarial outliers, even though the first five local testing procedures considered become ineffective in this scenario. For additional experiments that offer a more detailed view of the performance of our empirical approximation of the LMPI local testing procedure in different settings, please refer to Figures A6A7 in Appendix A4.1.
Figure 4: Performance of ACODE on synthetic data with adverserially hidden outliers exhibiting underdispersed conformity scores. Most local testing procedures cannot detect these outliers, but a data-driven approximation of the LMPI test enables our method to achieve high power. Other details are as in Figure 2.
@@ 3.2 Experiments with High-Energy Particle Collision Data
In each experiment, we randomly select a test set of 10,000 events, among which the proportion of outliers is a control parameter varied between 0 and 0.15. We also randomly sample three disjoint subsets of inliers, to be used for training (cardinality 10,000), calibration (cardinality 1000), and tuning (cardinality 1000), respectively. ACODE is then applied as described in Section 3.1, leveraging the same suites of 6 classification algorithms and 5 local testing procedures. The additional local testing procedure described in Section 3.1.4 (the data-driven approximation of the LMPI oracle) is omitted here, for simplicity. This omission is due to its somewhat higher computational cost of its closed testing shortcut and the lack of clear advantages observed with these data.
@@
Overall, ACODE yields more informative inferences compared to individual-level outlier identification, especially when leveraging the WMW test or Fisher’s combination method. Further, ACODE achieves near-oracle performance with respect to the selections of classifier and local testing procedure, without incurring in selection bias.
Figure 5 describes related results from experiments in which ACODE is applied to construct a 90% lower confidence bound for the number of outliers in a data-driven subset of test points, selected as those with the largest conformity scores. Similar to Figure 3, in these experiments the scores are computed by a single classifier, AdaBoost, to facilitates the interpretation of the findings. The results confirm the LMPI test for Lehmann’s alternative with $k=3$ tends to lead to more informative lower bounds compared to other local testing procedures, including the WMW test. Moreover, the adaptive version of ACODE again approximately maximizes power by selecting the testing procedure automatically. Additional results, with qualitatively similar conclusions, are presented in Appendix A4.2.
Figure 5: Median values for a 90% lower confidence bound on the number of outliers within an adaptively selected subset of 1000 test points, in numerical experiments otherwise similar to those of Figure 2. The results are shown as a function of the proportion of selected test points and of the total number of outliers in the test set. The dashed curve corresponds to the true number of outliers in this selected set.
@@ 4 Discussion
Lastly, an intriguing question is whether the optimality of local tests extends to closed testing procedures, a concept intertwined with the notion of admissibility. Although this issue remains unexplored, a starting point may be found in Goeman et al. (2021), which provides a general discussion on admissibility and its implications for closed testing.
Software Availability
An open-source software implementation of the methods described in this paper, and code needed to reproduce all results, are available online at https://github.com/msesia/conformal_nout.
@@
WMW Permutations Exact Rare/Weak effects Wilcoxon (1945), Mann and Whitney (1947)
@@
A permutation test is exactly valid, i.e. $\mathbb{P}[\phi^{\mathrm{Perm}}_{S}=1\mid H_{S}]=\alpha$, for any choice of test statistic under the sole assumption that the scores $(X_{1},\ldots,X_{m},(Y_{i})_{i\in I_{0}})$ are exchangeable. However, achieving exact validity for all values of $\alpha$ is generally not feasible due to the discreteness of the permutation distribution. Nevertheless, it is possible to ensure exactness through a mathematical artifice using a randomized critical region.
@@
The precise computation of $\alpha^{\mathrm{Perm}}$ for the Simes statistic requires (${n+|S|\choose|S|}$ permutations, representing the number of ways, disregarding order, that $|S|$ test units can be chosen from among $n+|S|$ units. However, this $\alpha^{\mathrm{Perm}}$ does not depend on the scores $(W_{1},\ldots,W_{m+|S|})$ but only on $m$, $|S|$ and $\alpha$, making it possible to tabulate. This is because the permutation-based Simes test is a rank test.
@@
$\frac{|S|}{m+1}\left\lfloor\alpha\frac{m+1}{|S|}\right\rfloor\leq\mathbb{P}[\phi^{\mathrm{Simes}}_{S}=1\mid H_{S}]\leq\mathbb{P}[\phi^{\mathrm{SimesPerm}}_{S}=1\mid H_{S}]\leq\alpha,$ (A14)
@@
If $\alpha^{\mathrm{Perm}}<\alpha$, then $\phi^{\mathrm{SimesPerm}}_{S}\leq\phi^{\mathrm{Simes}}_{S}$. We need to show that if $\alpha^{\mathrm{Perm}}<\alpha$, then $\phi^{\mathrm{SimesPerm}}_{S}=\phi^{\mathrm{Perm}}_{S}$ for all $(p_{j})_{j\in S}$. We will derive a contradiction. Suppose $\phi^{\mathrm{Simes}}_{S}>\phi^{\mathrm{PermSimes}}_{S}$ for some $(\tilde{p}_{j})_{j\in S}$. The corresponding test statistic $\tilde{T}^{\mathrm{Simes}}_{S}$ must be $\alpha^{\mathrm{Perm}}<\tilde{T}^{\mathrm{Simes}}_{S}\leq\alpha$. Since $\tilde{T}^{\mathrm{Simes}}_{S}=T_{\tilde{\pi}}$ for some $\tilde{\pi}\in\Pi$, it follows that $|\Pi|^{-1}\sum_{\pi\in\Pi}\mathds{1}\{T_{\pi}\leq\alpha\}>\alpha$ and we have a contradiction because $\mathbb{P}[\phi^{\mathrm{Simes}}_{S}=1\mid H_{S}]\leq\alpha$.
@@
An important class of permutation tests is defined in terms of sample ranks, known as rank tests. These tests rely on the ranks of the data, making them invariant to monotone transformations, as these transformations don’t alter the relative ranking of observations.
For any $S\subseteq[n]$ and any $j\in S$, let $R_{j}$ denote the rank of $Y_{j}$ among $(X_{1},\ldots,X_{m},(Y_{j})_{j\in S})$. The dependence of $R_{j}$ on $S$ is left implicit to simplify the notation, since this does not create ambiguity.
@@
When $S=\{j\}$, the WMW two-sample test, with the 1st sample being $X_{1},\ldots,X_{n}$ and the 2nd “sample” being just $Y_{j}$, rejects $H_{j}$ if the rank of $Y_{j}$ in the sequence $(X_{1},\ldots,X_{n},Y_{j})$ is large or, equivalently, if the conformal p-value $p_{j}$ in  (3) is small. Therefore, the WMW test provides a particularly intuitive bridge between the modern framework of conformal inference and the classical world of two-sample rank tests (Kuchibhotla, 2020).

### v3

@@ 2.6 Data-Driven Tuning
We extend Algorithm 2 to leverage a potentially diverse suite of classifiers and local testing procedures, with the goal of approximately maximizing power while controlling type-I error. Our solution, outlined in Algorithm 3, is inspired by approaches proposed by Liang et al. (2024) and Marandon et al. (2024) in the context of individual outlier identification under FDR control. The key idea is to embed a data-driven tuning step within an additional layer of sample splitting. As detailed below, this extra sample splitting is essential to preserve the validity of local tests after selecting the most effective classifier and local testing procedure; at the same time, provided the available sample size is sufficiently large, it typically does not lead to a substantial loss of power relative to data-driven selection without splitting (an intuitive but theoretically invalid benchmark).
@@
Test data $D^{\text{test}}=\{Z_{m+1},\ldots,Z_{m+n}\}$, Significance level $\alpha\in(0,1)$.
@@
6      Compute $d_{k,l}(R)$ and $\hat{p}_{k,l}(R)$ by applying Algorithm 2 based on $\mathcal{A}^{(k)}$ and $\phi^{(l)}$, using $\mathcal{D}^{\text{train}}$ for training, $\mathcal{D}^{\text{tune}}$ for calibration, and $\tilde{\mathcal{D}}^{\text{test}}$ as the test set.
7
@@
9Find $(\hat{k}(R),\hat{l}(R))=\arg\max_{k\in[K],l\in[L]}d_{k,l}(R)$; if there are multiple maximizers, select the one with the smallest $p$-value $\hat{p}_{k,l}(R)$.
10
11Query $S\subseteq[n]$: compute $d(S)$ and $\hat{p}(S)$ with Algorithm 2 based on $\mathcal{A}^{(\hat{k}(R))}$ and $\phi^{(\hat{l}(R))}$, using $\mathcal{D}^{\text{train}}\cup\mathcal{D}^{\text{tune}}$ for training, $\mathcal{D}^{\text{cal}}$ for calibration, and $\mathcal{D}^{\text{test}}$ as test set.
Output : A $(1-\alpha)$ lower bound $d(S)$ for the number of outliers in $S$, and a $p$-value $\hat{p}(S)$ for testing $H_{S}$, for any queried subset $S\subseteq[n]$.
12
@@
The tuning module of Algorithm 3 sees the calibration and test data only through the lenses of the unordered collection $\{Z_{1},\ldots,Z_{m}\}\cup\{Z_{m+1},\ldots,Z_{m+n}\}$. Consequently, when Algorithm 2 is applied again in the inference step of Algorithm 3, using $(Z_{1},\ldots,Z_{m})$ as a calibration set and $(Z_{m+1},\ldots,Z_{m+n})$ as a test set, the inlier scores within $(Z_{1},\ldots,Z_{m+n})$ are still exchangeable conditional on the selected classifier and testing procedure.
@@
Algorithm 3 changes this picture a little. As we condition on the selected classifier and closed testing procedure, the scores of the inliers among $(Z_{1},\ldots,Z_{m+n})$, although still exchangeable, can no longer be independent, irrespective of whether they were initially obtained through one-class or binary classification.
Relaxing the i.i.d. assumption on the inliers to exchangeability does not affect inference. Theorems 1 and A2 show that the large-sample approximations of the null distributions of Shiraishi’s statistic and its adaptive version remain valid under exchangeability. Moreover, for the other local testing procedures considered in this paper, namely Fisher’s and Simes’ combinations, the joint distribution of null conformal $p$-values is unchanged whether the inliers are i.i.d. or exchangeable, as shown in Gazin et al. (2024, Proposition 2.2). Formally, Algorithm 3 outputs a simultaneous lower bound $d(S)$ that satisfies (11), provided that exchangeability among the inliers holds. Thus, Algorithm 3 can often achieve high power without selection bias, as previewed in Figure 1 and confirmed in Section 3.
Finally, Algorithm 3 requires specification of a subset of interest $R$, with default choice $R=[n]$. ACODE then selects the pair of classifier and local test yielding the largest lower bound $d(R)$ on the number of outliers in $R$. In principle, multiple choices of $R$ may be considered, and the final selection of the (classifier, local test) pair may be made post hoc by comparing their results, provided that this choice is made before entering line 7 of the Algorithm. If the goal is instead to optimize rejection of the global null hypothesis (detection), rather than to maximize the lower bound (enumeration), the algorithm should select the (classifier, local test) pair that yields the smallest $p$-value. That is, $(\hat{k}(R),\hat{l}(R))=\arg\min_{k\in[K],l\in[L]}\hat{p}_{k,l}(R)$ with $R=[n]$.
@@ 3 Empirical Demonstrations
Section 3.1 applies ACODE to synthetic data, while Section 3.2 considers the 2020 LHCO dataset (Kasieczka et al., 2021). Additional experiments are reported in Appendix A6. In particular, Appendix A6.3 presents results on datasets previously studied by Bates et al. (2023) and Marandon et al. (2024). Consistent with Marandon et al. (2024), binary classifiers achieve higher power on those relatively low-dimensional data. Yet, as known from Liang et al. (2024) and confirmed in this section, one-class classifiers can yield higher power on different data sets, highlighting the practical benefit of ACODE’s adaptability.
@@ 3.1.1 Setup
We apply ACODE on synthetic data from a distribution inspired by Liang et al. (2024) and Bates et al. (2023). Each observation $Z_{i}\in\mathbb{R}^{1000}$ is sampled from a multivariate Gaussian mixture $P_{Z}^{a}$, such that $Z_{i}=\sqrt{a}\,V_{i}+W_{i}$, for some constant $a\geq 0$ and vectors $V_{i},W_{i}\in\mathbb{R}^{1000}$. The inliers correspond to $a=1$ and the outliers to $a=0.7$. The elements of $V_{i}$ are standard Gaussian, while each element of $W_{i}$ is independent and uniformly distributed on a discrete set $\mathcal{W}\subseteq\mathbb{R}^{1000}$ with $|\mathcal{W}|=1000$. The vectors in $\mathcal{W}$ are independently sampled from Uniform($3,3]^{1000}$) prior to the first experiment. The numbers of inliers in the training, calibration, and tuning sets are by default 1000, 750, and 250, respectively. As shown in Figure A2 (Appendix A6.1), ACODE’s performance is quite stable across a wide range of tuning splits, degrading only when tuning data are extremely scarce.
@@
The testing procedures leveraged by ACODE are: Simes test with and without Storey’s correction, Fisher’s combination method, the WMW test, and Shiraishi’s LMPR test from Section 2.3.2, applied using $G(F)=F^{3}$. This choice of $G$ is optimal against Lehmann’s alternative with $k=3$, which may or may not fit our data well. An implementation of the Shiraishi test based on a data-driven estimate of $G$ will be considered later.
@@ 3.1.3 Selective Outlier Enumeration
Figure 3 reports on related experiments in which the goal is to construct a 90% lower confidence bounds for the number of outliers within a data-driven subset of test points, selected as those with the largest conformity scores. To facilitate the interpretation of these experiments, ACODE is applied using a fixed classification algorithm, a one-class support vector machine. This ensures the test subsets are always selected based on comparable conformity scores in all repetitions of the experiments. We do not apply ACODE with the Shiraishi local test here due to its higher computational cost when the selected subset differs from the full test set; see Table A3. These results further demonstrate how the efficacy of different local testing procedures can vary in different situations. Simes’ method performs better with very small selected sets, while the WMW and Fisher’s test excel in cases with moderately large selected sets. Once more, ACODE can approximately maximize power by autonomously identifying the most effective local testing procedure for each scenario.
@@ 3.1.4 Hunting for Adversarial Anomalies
We now apply ACODE leveraging an additional local testing procedure, utilizing a more flexible implementation of the LMPR approach outlined in Section 2.3.2. This approach enables our method to even detect elusive anomalies hidden by an adversary.
@@
Without prior knowledge of the distribution of outlier scores, ACODE is applied as in previous experiments, using the same toolbox of six machine learning classifiers, but including a sixth local testing procedure. This procedure is a practical approximation of the Shiraishi test described in Section 2.3.2, using an empirical estimate $\hat{g}$ of $g$. This estimate is obtained by fitting a mixture of Beta distributions, while preserving the exchangeability between calibration and test scores, as detailed in Appendix A2.
The results reported in Figure 4 demonstrate that our method effectively detects these elusive adversarial outliers, even though the first five local testing procedures considered become ineffective in this scenario. For additional experiments that offer a more detailed view of the performance of our empirical approximation of the Shiraishi local testing procedure in different settings, please refer to Figures A11A12 in Appendix A6.1.
Figure 4: Empirical 90th percentile of the 90% lower confidence bound on the number of outliers in the test set, computed by ACODE on synthetic data with adversarially hidden outliers exhibiting underdispersed conformity scores. Most local testing procedures cannot detect these outliers, but a data-driven approximation of the Shiraishi test enables our method to achieve high power. Other details are as in Figure 2.
@@ 3.2 Experiments with High-Energy Particle Collision Data
In each experiment, we randomly select a test set of 10,000 events, among which the proportion of outliers is a control parameter varied between 0 and 0.15. We also randomly sample three disjoint subsets of inliers, to be used for training (cardinality 10,000), calibration (cardinality 1000), and tuning (cardinality 1000), respectively. ACODE is then applied as described in Section 3.1, leveraging the same suites of 6 classification algorithms and 5 local testing procedures. The additional local testing procedure described in Section 3.1.4 (the data-driven approximation of the Shiraishi oracle) is omitted here, for simplicity. This omission is due to its somewhat higher computational cost of its closed testing shortcut and the lack of clear advantages observed with these data.
@@
Overall, ACODE yields more informative inferences compared to individual-level outlier identification, especially when leveraging the WMW test or Fisher’s combination method. Further, ACODE achieves near-oracle performance with respect to the selections of classifier and local testing procedure, without incurring in selection bias. Additional results, with qualitatively similar conclusions, are presented in Appendix A6.2.
@@ 4 Discussion
In the future, one could investigate whether faster closed-testing shortcuts can be developed for the Shirashi local test, which tends to be powerful but is relatively computationally expensive as currently implemented, especially in the adaptive selection regime. A further extension is to integrate alternative procedures for estimating the number of outliers, such as the approach of Gazin et al. (2024), which can be interpreted as approximate shortcuts to closed testing and could be incorporated into ACODE, as discussed in Appendix A3.6. Lastly, an intriguing question is whether the optimality of local tests extends to closed testing procedures, a concept intertwined with the notion of admissibility. Although this issue remains unexplored, a starting point may be found in Goeman et al. (2021), which provides a general discussion on admissibility and its implications for closed testing.
AI Usage. ChatGPT-5.2 for language editing.
Funding. Co-funded by the European Union (ERC, BigBayesUQ, project number: 101041064). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@
WMW Permutations Exact Moderately Sparse/Weak effects Wilcoxon (1945), Mann and Whitney (1947)
A variant of the Simes’ test incorporating Storey’s estimator for the proportion of nulls has also been proven to be valid and may lead to improvement if the proportion of true nulls is not close to 1. See Appendix A1.1 for more details and the proof of the exactness of the Simes’ test when $\alpha(m+1)/|S|$ is an integer.
§ The asymptotic approximation proposed by Bates et al. (2023) works well even with moderate values of $m$ and $|S|$. Our empirical observations suggest that permutations can sometimes enhance power, as the asymptotic approximation tends to be somewhat conservative when both $m$ and $|S|$ are small. Therefore, for very small samples, one may want to embed Fisher’s combination statistic within a permutation test, which increases the computational cost but ensures exact finite-sample validity under minimal exchangeability assumptions. We refer to Section A1.3 for a more detailed review.
Permutation-based methods offer exact finite-sample validity as long as the inlier scores in $(X_{1},\ldots,X_{m},(Y_{j})_{j\in I_{0}})$ are exchangeable.
The critical value for the WMW test may be calculated either exactly, in small samples, or through accurate asymptotic approximations. See Appendix A1.4 for further details.
Table A1: Standard approaches for testing $H_{S}$.
@@
A permutation test is exactly valid, i.e. $\mathbb{P}[\phi^{\mathrm{Perm}}_{S}=1\;;\;H_{S}]=\alpha$, for any choice of test statistic under the sole assumption that the scores $(X_{1},\ldots,X_{m},(Y_{i})_{i\in I_{0}})$ are exchangeable. However, achieving exact validity for all values of $\alpha$ is generally not feasible due to the discreteness of the permutation distribution. Nevertheless, it is possible to ensure exactness through a mathematical artifice using a randomized critical region.
@@
The precise computation of $\alpha^{\mathrm{Perm}}$ for the Simes statistic requires $n+|S|\choose|S|$} permutations, representing the number of ways, disregarding order, that $|S|$ test units can be chosen from among $n+|S|$ units. However, this $\alpha^{\mathrm{Perm}}$ does not depend on the scores $(W_{1},\ldots,W_{m+|S|})$ but only on $m$, $|S|$ and $\alpha$, making it possible to tabulate. This is because the permutation-based Simes test is a rank test.
@@
$\frac{|S|}{m+1}\left\lfloor\alpha\frac{m+1}{|S|}\right\rfloor\leq\mathbb{P}[\phi^{\mathrm{Simes}}_{S}=1\;;\;H_{S}]\leq\mathbb{P}[\phi^{\mathrm{SimesPerm}}_{S}=1\;;\;H_{S}]\leq\alpha,$ (A16)
@@
If $\alpha^{\mathrm{Perm}}<\alpha$, then $\phi^{\mathrm{SimesPerm}}_{S}\leq\phi^{\mathrm{Simes}}_{S}$. We need to show that if $\alpha^{\mathrm{Perm}}<\alpha$, then $\phi^{\mathrm{SimesPerm}}_{S}=\phi^{\mathrm{Perm}}_{S}$ for all $(p_{j})_{j\in S}$. We will derive a contradiction. Suppose $\phi^{\mathrm{Simes}}_{S}>\phi^{\mathrm{PermSimes}}_{S}$ for some $(\tilde{p}_{j})_{j\in S}$. The corresponding test statistic $\tilde{T}^{\mathrm{Simes}}_{S}$ must be $\alpha^{\mathrm{Perm}}<\tilde{T}^{\mathrm{Simes}}_{S}\leq\alpha$. Since $\tilde{T}^{\mathrm{Simes}}_{S}=T_{\tilde{\pi}}$ for some $\tilde{\pi}\in\Pi$, it follows that $|\Pi|^{-1}\sum_{\pi\in\Pi}\mathds{1}\{T_{\pi}\leq\alpha\}>\alpha$ and we have a contradiction because $\mathbb{P}[\phi^{\mathrm{Simes}}_{S}=1\;;\;H_{S}]\leq\alpha$.
@@
An important class of permutation tests is defined in terms of sample ranks, known as rank tests. These tests rely on the ranks of the data, making them invariant to monotone transformations, as these transformations do not alter the relative ranking of observations.
For any $S\subseteq[n]$ with cardinality $s$, let $R_{1},\ldots,R_{m+s}$ denote the ranks of $(X_{1},\ldots,X_{m},(Y_{j})_{j\in S})$. The dependence of $R_{1},\ldots,R_{m+s}$ on $S$ is left implicit to simplify the notation, since this does not create ambiguity.
@@
When $S=\{j\}$, the WMW two-sample test, with the 1st sample being $X_{1},\ldots,X_{n}$ and the 2nd “sample” being just $Y_{j}$, rejects $H_{j}$ if the rank of $Y_{j}$ in the sequence $(X_{1},\ldots,X_{n},Y_{j})$ is large or, equivalently, if the conformal $p$-value $p_{j}$ in (3) is small. Therefore, the WMW test provides a particularly intuitive bridge between the modern framework of conformal inference and the classical world of two-sample rank tests (Kuchibhotla, 2021).